The distribution of the number of months in service for the fleet of cars is bell-shaped and has a mean of 45 months and a standard deviation of 9 months


A company has a policy of retiring company cars, this policy looks at number of miles driven, purpose of trips, style of car and other features. The distribution of the number of months in service for the fleet of cars is bell-shaped and has a mean of 45 months and a standard deviation of 9 months. Using the empirical rule (as presented in the book), what is the approximate percentage of cars that remain in service between 54 and 63 months?  
Answer #1

The empirical rule for normal distribution is as follows:

µ = 45

σ = 9

µ + σ = 45+9 = 54

µ + 2σ = 45+2*9 = 63

From the graph, it can be seen that percentage between µ + σ and µ + 2σ is 13.5%

So, answer is: 13.5 %